# ZMP-HH/24-01 The Standard Model from String Theory: What Have We Learned?

### 03 Jan 2024 v1

@@ 3.2 Gauge Coupling Unification
Gauge coupling unification is most natural if the three SM gauge group factors share a common group theoretic origin in the sense of a Grand Unified Theory (GUT). A key challenge in GUT model building is then the breaking of the GUT group to the SM gauge group . In purely field theoretic constructions, this is achieved by a Higgs mechanism. Apart from the question of how to generate the potential of a GUT Higgs dynamically, this approach is inflicted with phenomenological challenges such as the notorious doublet-triplet splitting problem in GUTs. It is therefore an interesting question for model building if string theory provides new mechanisms to address these challenges in a GUT context. This question comes in two parts: First, which types of GUT groups appear naturally, and second, which mechanisms for GUT breaking are available in addition to the field theoretic Higgs mechanism? Besides naturally accomodating phenomenologically appealing GUT groups, stringy GUTs offer a topological mechanism to realize the GUT breaking, which is intimately tied to the higher-dimensional nature of the theory. Namely, such constructions admit background values for suitable components of the gauge fields along the extra dimensions, as already mentioned in Section 2.1.
@@
These general ideas are very concretely realized already in the perturbative heterotic string, in which a gauge background on the compactification space reduces one of the ten-dimensional gauge factors to a GUT group such as , Spin or . The GUT group can then be further broken topologically to by extra discrete gauge backgrounds. In the context of compactifications on smooth Calabi-Yau threefolds this last step requires discrete Wilson lines, as mentioned already in Section 2.1. This approach has resulted in the exact SM spectrum embedded into a GUT scenario Bouchard:2005ag (43, 44), including the possibility of vector-like Higgs pairs.
@@
Alternatively, the small parametric separation between the GUT scale, GeV, and the four-dimensional Planck scale GeV, can be viewed as motivation for a local approach to unification. The starting point is a gauge sector localized on a spacetime filling brane whose internal directions extend along a shrinkable cycle on the compactification manifold. The embedding of the Standard Model into a GUT furthermore turns out to be particularly natural in a framework that admits exceptional gauge algebras. Taken together, these two considerations – localization of gauge degrees of freedom on branes and appearance of exceptional gauge degrees of freedom – point to the idea of F-theory GUTs initiated in BeasleyHeckmanVafaI (27, 28, 29, 30) (see Weigand:2010wm (80, 25) for more detailed reviews).
@@
Example: SU(5) F-theory GUTs Particular attention has been devoted to the GUT scenario in which an gauge group is engineered on stack of 7-branes. These wrap a holomorphic four-cycle on the compactification space , which in F-theory is identified with the base of an elliptically fibered Calabi-Yau four-fold . Within a local approach should be a shrinkable divisor and hence a generalized del Pezzo surface even though this restriction is not strictly required.
@@
Despite the increasing tension with LHC results, let us focus for simplicity on an supersymmetric GUT scenario. Apart from the gauge degrees of freedom propagating along , there must exist three generations of GUT matter in the representations and of SU(5). The SM Higgs sits in a vector-like pair of . All these are localized on curves where intersects additional 7-branes. In a purely local approach, these additional branes are not specified as they extend away from the localized GUT sector. Part of explaining the coupling to gravity is about specifying these further non-GUT branes.
@@
If the internal space along which the gauge field propagates admits non-trivial one-cycles, there can be topologically non-trivial Wilson line backgrounds, similar to the aforementioned GUT breaking in heterotic compactifications on non-simply Calabi-Yau threefolds. Even in absence of such one-cycles, the background gauge field can acquire quantized background values. Geometrically such configurations correspond to non-zero field strength VEVs along two-cycles. Specifically, the breaking is implemented by a non-trivial VEV for the hypercharge field strength along , .
Typically non-trivial flux backgrounds lead to a Green-Schwarz mechanism, which in the present context would make the massive. This unacceptable effect can also be circumvented thanks to the brane-localization of the GUT gauge group in the compact extra dimensions: As long as the class of on is topologically trivial within , no Green-Schwarz mass is induced Buican:2006sn (72, 81, 29). Along the way, this GUT breaking mechanism in principle allows for topological ways to solve the notorious doublet-triplet splitting problem. A recent overview of some of the proposed constructions can be found in Marchesano:2022qbx (5) along with open questions related to the embedding of the particle sector into a full theory including gravity.
@@ 3.3 Mechanisms to Generate Hierarchies
Generating hierarchies between scales or couplings is one of the key challenges in particle physics. In string theory, all scales (measured in units of the string scale ) and couplings are dynamical in the sense that they are determined by vacuum expectation values of scalar fields. This paves the way for new mechanisms to generate hierarchies.
We will discuss two examples for the generation of such hierarchies which cannot be explained in purely four-dimensional quantum field theory. The first relies on extra dimensions, the second on non-perturbative gravitational dynamics. While both effects could in principle be discussed independently of string theory, the latter naturally accommodates them and makes them computable in UV-complete frameworks.
@@ 3.3.1 Large extra dimensions and warpings
The probably most prominent mechanism to generate hierarchies in scales relies on the size of the extra dimensions Arkani_Hamed_1998 (82, 83) and the possibility of a warp factor Randall_1999 (84).
@@
where , , and , , are coordinates along the visible and extra dimensions, respectively. The warp factor depends only on the internal coordinates. Non-trivial warp factors naturally occur in flux compactifications or as a consequence of brane backreactions Verlinde_2000 (85, 86, 87).
Just like large extra dimension, warping has the potential to create a hierarchy between ten-dimensional Planck scale and the effective Planck scale in four dimensions. These are related by
@@ 3.3.2 Selection rules and non-perturbative effects
In many phenomenological scenarios selection rules constrain the set of possible operators in the effective action. In the low-energy theory these selection rules may appear as global symmetries. In quantum gravity, however, no exact global symmetries are believed to exist Banks:1988yz (89, 90) and hence such symmetries must either be gauge symmetries at high energies or be broken. Symmetries broken only by subleading effects give rise to approximate selection rules, and explain why some operators in the effective theory are strongly suppressed.
Consider an operator with charge under what appears to be a global symmetry at the level of the perturbative low energy effective action. Here perturbative may refer either to the expansion in the string coupling or to the worldsheet expansion. In string theory, such s are always associated with dynamical abelian gauge fields which acquire a Stückelberg mass term . Above this scale they behave as gauge symmetries while at energies below they appear as perturbative global selection rules which prevent any operator of charge from appearing in the perturbative effective action. However, an inevitable consequence of the Stückelberg mechanism is that non-perturbative effects break this perturbative selection rule to a discrete subgroup or even completely. This effect relies on the existence of instantons carrying an effective charge under the Blumenhagen:2006xt (91, 92, 93, 94). An instanton of charge can give rise to an operator
in the effective action. This leads to non-perturbatively suppressed terms in the effective action and hence serves as a mechanism to naturally generate hierarchically small couplings.
@@
where is the volume in string units. Integrating the Ramond-Ramond form over gives rise to an axion. Precisely this axion participates in the Stückelberg mechanism making the massive and hence transforms under the symmetry as . This results in the charge of the instanton operator . The mechanism is generic in string theory and not restricted to Type II constructions. For instance, in heterotic compactifications, stringy instantons are either worldsheet instantons (and hence non-perturbative in rather than ) or NS5-brane instantons (which are non-perturbative in ).
@@
In the presence of an instanton of charge , the abelian selection rule is broken completely at the non-perturbative level. By contrast, if all instantons carry charges which are multiples of some integer value , then only operators of charge are generated. The global symmetry (in the sense explained above) is broken to a symmetry, which is, by definition, exact at the non-perturbative level. Note that in the fully-fledged string (or quantum gravity) theory, the symmetry is really a discrete gauge symmetry Berasaluce-Gonzalez:2011gos (97, 98), rather than a global symmetry, in agreement with the lore that all exact symmetries must be gauged in quantum gravity Banks:2010zn (90). This way of generating a symmetry is qualitatively different from the familiar field theoretic mechanism of Higgsing a symmetry by a Higgs field of charge : In this case, the symmetry is restored at finite distance in the moduli space (corresponding to vanishing Higgs field), while here the is restored only at infinite distance in field space, where .
@@
An intrinsically string theoretic type of discrete symmetries are modular symmetries. Combined with conventional discrete symmetries these offer interesting avenues to understand the flavor hierarchies in the SM, as reviewed in Nilles:2023shk (100).
The key lesson from this discussion is two-fold: First, all questions related to global symmetries and selection rules must be analyzed in the context of a quantum gravity theory, rather than for the particle sector in isolation. Second, gravitational dynamics such as charged gravitational instantons and their concrete string theoretic realization as charged stringy instantons offer a way to create hierarchical couplings in effective theories.
@@ 4.1 Neutrino Sector
Right-handed neutrinos are not part of the Standard Model, but they are probably the simplest explanation for the observed spectrum of neutrino masses and mixing angles, via the see-saw mechanism. They oftentimes appear in semi-realistic string theory constructions, either because the model is based on a GUT gauge group like Spin or , or because it contains the perturbative symmetry , which becomes massive through a Stückelberg mechanism as described in Section 2.1.
The question is whether string compactifications can generate the necessary couplings that reproduce the see-saw mechanism. The potential difficulty is that, while right-handed neutrinos are not charged under the SM gauge group, they are charged under additional symmetries like , which forbid neutrino Majorana masses at the perturbative level. Nevertheless, two main mechanisms have been observed that overcome this obstruction. The first one has been observed in the context of heterotic orbifolds, see Faraggi:1993zh (101, 102, 103), and it is based on the existence of superpotential couplings of the form
where are the right-handed neutrinos, is either the compactification or the string scale and are some SM singlet fields which acquire large vevs upon moduli stabilization.
The second mechanism arises more naturally in Type II orientifold models Blumenhagen:2006xt (91, 92, 94), and it is based on non-perturbative effects breaking the symmetries that forbid Majorana masses, generating couplings of the form
@@
In both of these setups the right-handed neutrinos clearly belong to the gauge sector of the theory, as they are highly localized in the extra dimension and charged under some extension of the SM gauge group. This does not need to be the case in general, as has been pointed out in the context of F-theory GUTs Tatar:2009jk (104). It this setting the right-handed neutrino candidates are complex structure moduli of the bulk, which gain a mass via a moduli stabilization mechanism. Finally, in models without Majorana masses for right-handed neutrinos, there exist alternative mechanisms to generate realistic neutrino masses. In particular, instanton effects can also generate a superpotential term of the form
@@ 4.2 Moduli and Axions
A generic property of compactifications of string theory is the appearance of a plethora of light scalar fields whose masses lie at or below the compactification scale. In addition to the dilaton present already in ten dimensions, closed string moduli arise as massless fluctuations of the metric reduced along the compactified extra six dimensions. They are in one-to-one correspondence with flat deformations of the metric. Their vacuum expectation values parametrize the size and shape of the compactification space. For example, in Type II compactifications on Calabi-Yau threefolds, the closed string moduli space factors into deformations of the Kähler and the complex structure. Similarly, open string moduli are associated with deformations of spacetime-filling branes along the compactification space. In 4d compactifications with at most supersymmetry, various effects can induce a scalar potential on the moduli space, i.e. on the space of scalar fields, which may lead to non-vanishing mass terms for some or even all moduli in the minimum of the potential. Finding minima with no massless moduli is the goal of the moduli stabilization program as reviewed e.g. in Gra_a_2006 (17). The best studied examples include Type IIA compactifications with background fluxes stabilising both Kähler and complex structure moduli DeWolfe:2005uu (105) and Type IIB compactifications with background fluxes to stablize the complex structure moduli and quantum effects to stabilize the Kähler moduli Kachru:2003aw (106, 107). Similar effects also give rise to moduli masses of other classes of compactifications, see Section 4.3. Since the value of the moduli determine the couplings of the gauge sector, detailed questions about the interactions cannot be analyzed in isolation from the moduli problem.
@@
At least in scenarios with low-energy supersymmetry, one or several moduli tend to acquire masses around the supersymmetry breaking scale and hence are rather light. The associated challenges for cosmology are known as the cosmological moduli problem, see for instance Quevedo:2014xia (74) and references therein.
@@
Another generic property of string compactifications is the appearance of axions Svrcek:2006yi (108), at least prior to taking into consideration the effects of moduli stabilization. Axions arise naturally by reducing the Ramond-Ramond -form gauge potentials or the Neveu-Schwarz 2-form potential over appropriate cycles, for instance as , where is a -cycle on the compactification space. Open string axions arise from 1-form gauge potentials on D-branes, reduced over internal 1-cycles of the brane. In the low-energy supergravity theory, axions pair up with the moduli to form complex scalar fields. Unlike their saxionic partners, the geometric moduli, they enjoy a perturbative shift symmetry which is inherited from the gauge symmetry of the underlying -form potential. The perturbative shift symmetry can be broken by non-perturbative effects, i.e. by worldsheet or D-brane instantons in string theory, or in the presence of background sources such as gauge fluxes or wrapped branes. The axionic couplings in the EFT arise by dimensional reduction of the Chern-Simons couplings in a higher-dimensional effective action. For example, a D-brane has a Chern-Simons action of the form
@@
Analogous couplings appear in the heterotic string by dimensionally reducing the anomalous Green-Schwarz couplings. This may in principle give rise to an axion coupling to the strong sector as envisaged in the Peccei-Quinn solution to the strong CP problem. However, the phenomenological viability of the axion depends on its mass and axion decay constant, as reviewed for instance in Cicoli:2013ana (109). In a supersymmetric Minkowksi minimum, if existent at all Palti:2020qlc (110), the stabilization mechanism required to remove the moduli fields from the massless spectrum gives a mass of the same size to its axionic partner. Approximately massless axions are therefore expected to arise rather in AdS compactifications or in non-supersymmetric minima. In the spirit of Arvanitaki:2009fg (111), however, one can read the appearance of a QCD axion solving the strong CP problem as a constraint on a string compactification, and assuming its existence suggests the appearance of several, if not hundreds of axionic particles with interesting phenomenological properties. The reason is that a typical compactification manifold exhibits compact cycles over which the higher-dimensional -form potentials can be reduced.

### 08 Feb 2024 v2

@@ 3.2 Gauge Coupling Unification
Gauge coupling unification is most natural if the three SM gauge group factors share a common group theoretic origin in the sense of a Grand Unified Theory (GUT). A key challenge in GUT model building is then the breaking of the GUT group to the SM gauge group . In purely field theoretic constructions, this is achieved by a Higgs mechanism. Apart from the question of how to generate the potential of a GUT Higgs dynamically, this approach is inflicted with phenomenological challenges such as the notorious doublet-triplet splitting problem in GUTs. It is therefore an interesting question for model building if string theory provides new mechanisms to address these challenges in a GUT context. This question comes in two parts: First, which types of GUT groups appear naturally, and second, which mechanisms for GUT breaking are available in addition to the field theoretic Higgs mechanism? Besides naturally accomodating phenomenologically appealing GUT groups, stringy GUTs offer a topological mechanism to realize GUT breaking, which is intimately tied to the higher-dimensional nature of the theory. Namely, such constructions admit background values for suitable components of the gauge fields along the extra dimensions, as already mentioned in Section 2.1.
@@
These general ideas are very concretely realized already in the perturbative heterotic string. A detailed overview is given in Dienes:1996du (81). Focusing for concreteness on compactifications on a smooth space , a gauge background on reduces one of the ten-dimensional gauge factors to a GUT group such as , Spin or . The GUT group can then be further broken topologically to by extra discrete gauge backgrounds. In the context of compactifications on smooth Calabi-Yau threefolds this last step requires discrete Wilson lines, as mentioned already in Section 2.1. This approach has resulted in the exact SM spectrum embedded into a GUT scenario Bouchard:2005ag (43, 44), including the possibility of vector-like Higgs pairs.
@@
Alternatively, the small parametric separation between the GUT scale, GeV, and the four-dimensional Planck scale GeV, can be viewed as motivation for a local approach to unification. The starting point is a gauge sector localized on a spacetime filling brane whose internal directions extend along a shrinkable cycle on the compactification manifold. The embedding of the Standard Model into a GUT furthermore turns out to be particularly natural in a framework that admits exceptional gauge algebras. Taken together, both features – localization of gauge degrees of freedom on branes and appearance of exceptional gauge degrees of freedom – point to the idea of F-theory GUTs initiated in BeasleyHeckmanVafaI (27, 28, 29, 30) (see Weigand:2010wm (82, 25) for more detailed reviews).
@@
Example: SU(5) F-theory GUTs Particular attention has been devoted to the GUT scenario in which an gauge group is engineered on stack of 7-branes.333For constructions based on gauge groups Spin, or we refer to the discussion and references in Li:2023dya (83). These wrap a holomorphic four-cycle on the compactification space , which in F-theory is identified with the base of an elliptically fibered Calabi-Yau four-fold . Within a local approach should be a shrinkable divisor and hence a generalized del Pezzo surface even though this restriction is not strictly required.
@@
Despite the increasing tension with LHC results, let us focus for simplicity on an supersymmetric GUT scenario. Apart from the gauge degrees of freedom propagating along , there must exist three generations of GUT matter in the representations and of SU(5). The SM Higgs sits in a vector-like pair of . All these are localized on curves where intersects additional 7-branes. In a purely local approach, these additional branes are not specified as they extend away from the localized GUT sector. Part of explaining the coupling to gravity (to address the physics questions sensitive to global details of the model, cf. Section 3.1) is about specifying these further non-GUT branes.
@@
If the internal space along which the gauge field propagates admits non-trivial one-cycles, there can be topologically non-trivial Wilson line backgrounds, similar to the aforementioned GUT breaking in heterotic compactifications on non-simply connected Calabi-Yau threefolds. Even in absence of such one-cycles, the background gauge field can acquire quantized background values. Geometrically such configurations correspond to non-zero field strength VEVs along two-cycles. Specifically, the breaking is implemented by a non-trivial VEV for the hypercharge field strength along , .
Typically non-trivial flux backgrounds lead to a Green-Schwarz mechanism, which in the present context would make the massive. This unacceptable effect can also be circumvented thanks to the brane-localization of the GUT gauge group in the compact extra dimensions: As long as the class of on is topologically trivial within , no Green-Schwarz mass is induced Buican:2006sn (73, 84, 29). Chiral three-generation models on smooth compact elliptic 4-folds with 4-form fluxes (see Braun:2011zm (85) and references therein) of this type were first constructed in Marsano:2011hv (86, 87, 88).
Along the way, this GUT breaking mechanism in principle allows for topological ways to solve the notorious doublet-triplet splitting problem of GUTs. A recent overview of some of the proposed constructions can be found in Krippendorf:2015kta (89, 5) along with open questions related to the embedding of the particle sector into a full theory including gravity.
@@ 3.3 Mechanisms to Generate Hierarchies
Generating hierarchies between scales or couplings is one of the key challenges in particle physics. In string theory, all scales measured in units of the intrinsic string scale , and all dimensionless couplings are dynamical in the sense that they are determined by vacuum expectation values of scalar fields. This paves the way for new mechanisms to generate hierarchies. We will discuss two examples for the generation of such hierarchies which cannot be explained in purely four-dimensional quantum field theory. The first relies on extra dimensions, the second on non-perturbative gravitational dynamics. While both effects could in principle be discussed independently of string theory, the latter naturally accommodates them and makes them computable in UV-complete frameworks.
@@ 3.3.1 Large extra dimensions and warpings
The probably most prominent mechanism to generate hierarchies in scales relies on the size of the extra dimensions Arkani_Hamed_1998 (90, 91, 92, 93) and the possibility of a warp factor Randall_1999 (94).
@@
where , , and , , are coordinates along the visible and extra dimensions, respectively. The warp factor depends only on the internal coordinates. Non-trivial warp factors naturally occur in flux compactifications or as a consequence of brane backreactions Verlinde_2000 (95, 96, 97).
Just like large extra dimensions, warping has the potential to create a hierarchy between ten-dimensional Planck scale and the four-dimensional Planck scale. These are related by
@@ 3.3.2 Selection rules and non-perturbative effects
In many phenomenological scenarios selection rules constrain the set of possible operators in the effective action. In the low-energy theory these selection rules may appear as global symmetries. In quantum gravity, however, no exact global symmetries are believed to exist Banks:1988yz (99, 100) and hence such symmetries must either be gauge symmetries at high energies or be broken. Symmetries broken only by subleading effects give rise to approximate selection rules, and explain why some operators in the effective theory are strongly suppressed.
Consider an operator with charge under what appears to be a global symmetry at the level of the perturbative low energy effective action. Here perturbative may refer either to the expansion in the string coupling or to the worldsheet expansion. In string theory, such s are always associated with dynamical Abelian gauge fields which acquire a Stückelberg mass term . Above this scale they behave as gauge symmetries while at energies below they appear as perturbative global selection rules which prevent any operator of charge from appearing in the perturbative effective action. However, an inevitable consequence of the Stückelberg mechanism is that non-perturbative effects break this perturbative selection rule to a discrete subgroup or even completely. This effect results from the existence of instantons carrying an effective charge under the Blumenhagen:2006xt (101, 102, 103, 104). An instanton of charge can give rise to an operator with charge in the effective action,
This leads to non-perturbatively suppressed terms in the effective action and hence serves as a mechanism to naturally generate hierarchically small couplings. For explicit examples of the resulting suppression factors in the particle physics context we refer to the review Blumenhagen:2009qh (104) and references therein.
@@
where is the volume in string units. Integrating the Ramond-Ramond form over gives rise to an axion in the 4d EFT. Precisely this axion participates in the Stückelberg mechanism making the massive and hence transforms under the symmetry as . This results in the charge of the instanton operator . The mechanism is generic in string theory and not restricted to Type II constructions. For instance, in heterotic compactifications, stringy instantons are either worldsheet instantons (and hence non-perturbative in rather than in ) or NS5-brane instantons (which are non-perturbative in ).
@@
In the presence of an instanton of charge , the Abelian selection rule is broken completely at the non-perturbative level. By contrast, if all instantons carry charges which are multiples of some integer value , then only operators of charge are generated. The global symmetry (in the sense explained above) is broken to a symmetry, which is, by definition, exact at the non-perturbative level. Note that in the fully-fledged string (or quantum gravity) theory, the symmetry is really a discrete gauge symmetry Berasaluce-Gonzalez:2011gos (107, 108), rather than a global symmetry, in agreement with the lore that all exact symmetries must be gauged in quantum gravity Banks:2010zn (100). This way of generating a symmetry is qualitatively different from the familiar field theoretic mechanism of Higgsing a symmetry by a Higgs field of charge : In the latter case, the symmetry is restored at finite distance in the moduli space (corresponding to vanishing Higgs field), while here the is restored only at infinite distance in field space, where .
@@
An intrinsically string theoretic type of discrete symmetries are modular symmetries, e.g. with modular symmetry group . Combined with conventional discrete symmetries these offer interesting avenues to understand the flavor hierarchies in the SM, as reviewed in Nilles:2023shk (110).
The key lesson from this discussion is two-fold: First, all questions related to global symmetries and selection rules must be analyzed in the context of a quantum gravity theory, rather than for the particle sector in isolation. Second, gravitational dynamics such as charged gravitational instantons and their concrete string theoretic realization as charged stringy instantons offer a way to create hierarchical couplings in effective theories.Also, string theory provides a geometrical understanding of such mechanisms.
@@ 4.1 Neutrino Sector
Right-handed neutrinos are not part of the Standard Model, but they are probably the simplest explanation for the observed spectrum of neutrino masses and mixing angles, via the see-saw mechanism. They oftentimes appear in semi-realistic string theory constructions, either because the model is based on a GUT gauge group like Spin or that contains a right-handed neutrino in its basic representation, or because it contains the perturbative symmetry , which becomes massive through a Stückelberg mechanism as described in Section 2.1.
The question is whether string compactifications can generate the necessary couplings that reproduce the see-saw mechanism. This requires the generation of a Majorana neutrino mass. The potential difficulty is that, while right-handed neutrinos are not charged under the SM gauge group, they are charged under additional symmetries like , which forbid neutrino Majorana masses at the perturbative level. Nevertheless, two main mechanisms have been observed that overcome this obstruction. The first one has arisen in the context of heterotic orbifolds, see Faraggi:1993zh (111, 112, 113), which can include superpotential couplings of the form
Here are the right-handed neutrinos, is either the compactification or the string scale and are some SM singlet fields which acquire large vevs upon moduli stabilization.
The second mechanism arises more naturally in Type II orientifold models Blumenhagen:2006xt (101, 102, 104), and it is based on non-perturbative effects breaking the symmetries that forbid Majorana masses, generating couplings of the form
@@
In both of these setups the right-handed neutrinos clearly belong to the gauge sector of the theory, as they are highly localized in the extra dimension and charged under some extension of the SM gauge group. This does not need to be the case in general, as has been pointed out in the context of F-theory GUTs Bouchard:2009bu (114, 115). It this setting the right-handed neutrino candidates are complex structure moduli of the bulk Tatar:2009jk (115), which gain a mass via a moduli stabilization mechanism, or even Kaluza-Klein modes Bouchard:2009bu (114). Finally, in models without Majorana masses for right-handed neutrinos, there exist alternative mechanisms to generate realistic neutrino masses. In particular, instanton effects can also generate a superpotential term of the form
@@ 4.2 Moduli and Axions
A generic property of compactifications of string theory is the appearance of a plethora of light scalar fields whose masses lie at or below the compactification scale. In addition to the dilaton present already in ten dimensions, closed string moduli arise as massless fluctuations of the metric reduced along the compactified extra six dimensions. They are in one-to-one correspondence with flat deformations of the metric. Their vacuum expectation values parametrize the size and shape of the compactification space. For example, in Type II compactifications on Calabi-Yau threefolds, the closed string moduli space factors into deformations of the Kähler and the complex structure. Similarly, open string moduli are associated with deformations of spacetime-filling branes along the compactification space. In 4d compactifications with at most supersymmetry, various effects can induce a scalar potential on the moduli space, i.e. on the space of scalar fields, which may lead to non-vanishing mass terms for some or even all moduli in the minimum of the potential. Finding minima with no massless moduli is the goal of the moduli stabilization program as reviewed e.g. in Gra_a_2006 (17). The best studied examples include Type IIA compactifications with background fluxes stabilising both Kähler and complex structure moduli DeWolfe:2005uu (116) and Type IIB compactifications with background fluxes to stablize the complex structure moduli and quantum effects to stabilize the Kähler moduli Kachru:2003aw (117, 118). Similar effects also give rise to moduli masses of other classes of compactifications, see Section 4.3. Since the value of the moduli determine the couplings of the gauge sector, detailed questions about the size of the couplings cannot be analyzed in isolation from the moduli problem.
@@
At least in scenarios with low-energy supersymmetry, one or several moduli tend to acquire masses around the supersymmetry breaking scale and hence are rather light. This brings challenges for cosmology, known as the cosmological moduli problem,s see for instance Quevedo:2014xia (75) and references therein.
@@
Another generic property of string compactifications is the appearance of axions Svrcek:2006yi (119), at least prior to taking into consideration the effects of moduli stabilization. Axions arise naturally by reducing the Ramond-Ramond -form gauge potentials or the Neveu-Schwarz 2-form potential over appropriate cycles, for instance as , where is a -cycle on the compactification space. Open string axions arise from 1-form gauge potentials on D-branes, reduced over internal 1-cycles of the brane. In the low-energy supergravity theory, axions pair up with the moduli to form complex scalar fields. Unlike their partners, the geometric moduli, they enjoy a perturbative shift symmetry which is inherited from the gauge symmetry of the underlying -form potential. The perturbative shift symmetry can be broken by non-perturbative effects, i.e. by worldsheet or D-brane instantons in string theory, or in the presence of background sources such as gauge fluxes or wrapped branes. The axionic couplings in the EFT arise by dimensional reduction of the Chern-Simons couplings in a higher-dimensional effective action. For example, a D-brane has a Chern-Simons action of the form
@@
Analogous couplings appear in the heterotic string by dimensionally reducing the anomalous Green-Schwarz couplings. This may in principle give rise to an axion coupling to the strong sector as envisaged in the Peccei-Quinn solution to the strong CP problem. However, the phenomenological viability of the axion depends on its mass and axion decay constant, as reviewed for instance in Cicoli:2013ana (120). In a 4d supersymmetric Minkowksi minimum (if they exist at all Palti:2020qlc (121)), the stabilization mechanism required to remove the moduli fields from the massless spectrum gives a mass of the same size to its axionic partner. Approximately massless axions are therefore expected to arise rather in AdS compactifications or in non-supersymmetric minima. In the spirit of Arvanitaki:2009fg (122), however, one can read the appearance of a QCD axion solving the strong CP problem as a constraint on a string compactification, and assuming its existence suggests the appearance of several, if not hundreds of axionic particles with interesting phenomenological properties. The reason is that a typical compactification manifold exhibits compact cycles over which the higher-dimensional -form potentials can be reduced.